OpenStudy (anonymous):

Quadrilateral QRST is a square. If the measure of RS is 12x + 5 and QT is 8x + 15, find the value of QR how would i do this problem?

OpenStudy (anonymous):

12 x + 5 = 8x + 15 = QR

OpenStudy (anonymous):

4x = 10 x = 5/2 QR = 35

OpenStudy (anonymous):

how would i do this if i had the same problem but had to find the value of x?

OpenStudy (anonymous):

find x first then subsititute the value of x into the equation

OpenStudy (anonymous):

how do i find x?